# Binary mergers in the centers of galaxies: synergy between stellar flybys and tidal fields

### v1

@@ 7.2 Other astrophysical environments
We can see that for an average comet in the Oort Cloud we get $\mathcal{R}\simeq 9$ as opposed to the $\mathcal{R}\simeq 0.02$ obtained for compact object binaries in the GC. This implies that in the Oort Cloud flybys dominate over tidal fields, as the diffusion rate from these interactions is $\sim 9$ times faster than the torquing due to the tidal field. This is mainly due to the difference in the amount of flybys in a secular timescale, $\tau_{0}/t_{\mathrm{enc}}$. As shown in equation (A7), we get about 68 encounters for an average binary in the GC, while in the solar neighbourhood the average comet suffers $\sim 2\times 10^{4}$ flyby encounters (considering equation (26) as $\tau_{0}$ and $t_{\mathrm{enc}}\simeq 0.015\mathrm{Myr}$ as used by Heisler & Tremaine 1986). This increase in the amount of close encounters inevitably results in more diffusion of the angular momentum $\vec{j}$, hence the higher value of $\mathcal{R}$ obtained in this regime. Consistently, the numerical experiments by Heisler & Tremaine (1986) show the secular evolution is largely detuned due to the stellar flybys, making it hard to complete a clean secular cycle in the $\omega-e$ space (see figure 4 therein).
@@ 8 Conclusions
We would like to thank Antranik Sefilian, Carolina Charalambous, and Diego Muñoz for useful discussions. CP acknowledges support from CATA-Basal AFB-170002, ANID BASAL project FB210003, FONDECYT Regular grant 1210425, CASSACA grant CCJRF2105, and ANID+REC Convocatoria Nacional subvencion a la instalacion en la Academia convocatoria 2020 PAI77200076.
@@ A.3 Numerical experiments for $\mathcal{R}$
[…]

### v2

@@ 7.2 Other astrophysical environments
We can see that for an average comet in the Oort Cloud we get $\mathcal{R}\simeq 9$ as opposed to the $\mathcal{R}\simeq 0.02$ obtained for compact object binaries in the GC. This implies that in the Oort Cloud flybys dominate over tidal fields, as the diffusion rate from these interactions is $\sim 9$ times faster than the torquing due to the tidal field. This is mainly due to the difference in the amount of flybys in a secular timescale, $\tau_{0}/t_{\mathrm{enc}}$. As shown in equation (A7), we get about 68 encounters for an average binary in the GC, while in the solar neighbourhood the average comet suffers $\sim 2\times 10^{4}$ flyby encounters (considering equation (27) as $\tau_{0}$ and $t_{\mathrm{enc}}\simeq 0.015\mathrm{Myr}$ as used by Heisler & Tremaine 1986). This increase in the amount of close encounters inevitably results in more diffusion of the angular momentum $\vec{j}$, hence the higher value of $\mathcal{R}$ obtained in this regime. Consistently, the numerical experiments by Heisler & Tremaine (1986) show the secular evolution is largely detuned due to the stellar flybys, making it hard to complete a clean secular cycle in the $\omega-e$ space (see figure 4 therein). This in turn makes it difficult to achieve the high eccentricities required by a merger, as is discussed by Stegmann et al. (2024).
@@ 8 Conclusions
We would like to thank Antranik Sefilian, Carolina Charalambous, Diego Muñoz, Mark Dodici, and Scott Tremaine for useful discussions. We are also grateful to an anonymous referee for providing helpful comments on the maniscript. CP acknowledges support from CATA-Basal AFB-170002, ANID BASAL project FB210003, FONDECYT Regular grant 1210425, CASSACA grant CCJRF2105, and ANID+REC Convocatoria Nacional subvencion a la instalacion en la Academia convocatoria 2020 PAI77200076.
@@ A.3 Numerical experiments for $\mathcal{R}$
In order to confirm that the expression derived for $\mathcal{R}$ using the impulse approximation is accurate, we simulated a binary that evolves only by flyby interactions using REBOUND. From this numerical experiment, in which we considered $M_{\text{bin}}=10M_{\odot}$, $M_{p}=M_{\odot}$, $M_{\text{BH}}=4\times 10^{6}M_{\odot}$, $a_{\text{in}}=25$au, $a_{\text{out}}=0.3$pc and $v_{p}=200$kms-1, we calculated the mean displacement of $\|\vec{j}\|$ in different time windows until finding a value at which it converges after one secular timescale $\tau_{0}$. This is shown in Figure A2, where we can clearly see how $\mathcal{R}$ converges to $\approx 0.02$ after one secular timescale, which matches the value we retrieve using equation (22) when considering the same initial conditions. In order to keep this simulation as similar as possible to the derivations using the impulse approximation, we do not consider the changes in $a_{\text{in}}$ after each encounter, as this would change the amount of flybys per secular timescale.
@@
Appendix B Merger fraction and $j_{\text{crit}}$
To show that the optimal value of $\mathcal{R}$ at which the merger fraction peaks is proportional to the $j_{\text{crit}}$ chosen to classify mergers, a set of populations was simulated at different values for $\mathcal{R}$. Their merger fractions for three different values of $j_{\text{crit}}$ are plotted in figure B1 as a function of $\mathcal{R}$, and we can see that for smaller $j_{\text{crit}}$ the peak value tends towards a smaller $\mathcal{R}$ as well.
Figure B1: Merger fraction as defined by different values of $j_{\text{crit}}$ as a function of $\mathcal{R}$.